# Leave No Patient Behind: Enhancing Medication Recommendation for Rare Disease Patients

### v1

@@ Abstract.
Medication recommendation systems have gained significant attention in healthcare as a means of providing tailored and effective drug combinations based on patients’ clinical information. However, existing approaches often suffer from fairness issues, as recommendations tend to be more accurate for patients with common diseases compared to those with rare conditions. In this paper, we propose a novel model called Robust and Accurate REcommendations for Medication (RAREMed), which leverages the pretrain-finetune learning paradigm to enhance accuracy for rare diseases. RAREMed employs a transformer encoder with a unified input sequence approach to capture complex relationships among disease and procedure codes. Additionally, it introduces two self-supervised pre-training tasks, namely Sequence Matching Prediction (SMP) and Self Reconstruction (SR), to learn specialized medication needs and interrelations among clinical codes. Experimental results on two real-world datasets demonstrate that RAREMed provides accurate drug sets for both rare and common disease patients, thereby mitigating unfairness in medication recommendation systems.
Medication Recommendation, Fairness, Rare Disease
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USA.booktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAisbn: 979-8-4007-0431-4/24/07doi: 10.1145/XXXXXX.XXXXXXccs: Information systems Recommender systems
@@ 1. Introduction
Current approaches to medication recommendation primarily focus on maximizing overall accuracy by incorporating patients’ current medical conditions (Zhang et al., 2017; Wang et al., 2019) or clinical information from previous visits (Shang et al., 2019a, b; Yang et al., 2021b; Wu et al., 2022; Yang et al., 2021a; Sun et al., 2022). While significant progress has been made in comprehensive utilization these data, existing approaches encounter fairness issues in terms of recommendations across patient groups. This stems from the highly skewed distribution of clinical codes in Electronic Health Records (EHR) (Kim et al., 2019; Cowie et al., 2017) — a digital compilation of patients’ medical conditions, including diseases, procedures, medications and other clinical data during patient visits. As illustrated in Figure 1(a), a small subset of codes exhibit high prevalence, while the long tail comprises low-occurrence diseases in two real-world EHR datasets. Consequently, patients resided in the most common group receive significantly more accurate recommendations compared to those in the rarest group, as depicted in Figure 1(b). This disparity in accuracy across patient groups raises concerns regarding equity and fairness, ultimately undermining the reliability and utility of medication recommendation systems.
@@
  • Tailored medication needs [^1] We will further discuss the profile of rare disease patients in Appendix A.: The atypical pathophysiology of rare diseases often necessitates more precisely targeted pharmacological interventions. On the other hand, patients with rare diseases tend to have more complicated therapeutic regimens, warrant tailored medication recommendations.
@@ 2. Related Works
Medication Recommendation. Methods for medication recommendation can be broadly categorized into two groups: longitudinal and instance-based approaches. Longitudinal methods make use of patients’ longitudinal medical history, such as the approach taken by Choi et al. (Choi et al., 2016), who employ a two-level temporal attention mechanism. Similarly, Shang et al. (Shang et al., 2019a) pre-train their model on single-visit data [^3] Single-visit patients have no historical visits, i.e., have only one visit recorded in EHR. and fine-tune it using multi-visit data [^4] Multi-visit patients have at least two visits in EHR.. Additionally, Shang et al. (Shang et al., 2019b) incorporate a graph-augmented memory module and DDI graph to reduce adverse drug-drug interactions, while Yang et al. (Yang et al., 2021b) consider drug molecule structures for medication security. Wu et al. (2022) introduce a copy-or-predict mechanism, and Yang et al. (Yang et al., 2021a) propose a residual-based recurrent network. Sun et al. (Sun et al., 2022) present a causal model to address recommendation bias, and Yang et al. (Yang et al., 2023) leverage a molecular substructure-aware attentive method. More recently, Bhoi et al. (Bhoi et al., 2023) propose a fine-grained medication recommendation approach. However, it is important to note that many of these methods heavily rely on historical records, making them less suitable for single-visit patients (Choi et al., 2016; Shang et al., 2019a; Sun et al., 2022; Bhoi et al., 2023), or may exhibit reduced accuracy in such cases (Shang et al., 2019b; Yang et al., 2021b; Wu et al., 2022; Yang et al., 2021a, 2023). Instance-based methods focus on the current medical condition of patients. For instance, Zhang et al. (Zhang et al., 2017) treat the task as a multi-instance multi-label learning problem, utilizing an RNN model based on disease codes. Wang et al. (Wang et al., 2019) frame the task as an order-free Markov decision process, employing Deep Q Learning to capture correlations and adverse interactions between medicines. However, these models may not adequately capture patient-specific information, potentially leading to reduced accuracy. Despite the initial success of these methods, a common issue among them is fairness (Chen et al., 2023; Wang et al., 2023). Patients with common diseases tend to receive more accurate medication predictions compared to those with rare diseases. This discrepancy undermines the overall performance and impairs the reliability and usefulness of medication recommendation systems.
@@ 3. Problem Formulation
The record of patient $j$, denoted as $\mathcal{V}^{(j)}=\{\mathbf{d}^{(j)},\mathbf{p}^{(j)},\mathbf{m}^{(j)}\}$, comprises three components. Firstly, $\mathbf{d}^{(j)}=[d_{1},d_{2},\cdots,d_{m}]\in\mathcal{D}$ represents the sequence of diseases experienced by the patient, ordered by their priority or significance [^5] The sequence numbers of priority are provided in the EHR datasets (Johnson et al., 2016, 2023), which are labeled by qualified physicians., reflecting their relative importance in the patient’s healthcare journey. Similarly, $\mathbf{p}^{(j)}=[p_{1},p_{2},\cdots,p_{n}]\in\mathcal{P}$ captures the sequence of procedures, also ordered by priority. Here, $m$ and $n$ denote the lengths of the disease and procedure sequences, respectively. Lastly, $\mathbf{m}^{(j)}\in\{0,1\}^{|\mathcal{M}|}$ is a multi-hot vector indicating the medications prescribed to the patient. Each element in $\mathbf{m}^{(j)}$ corresponds to a specific medication in $\mathcal{M}$, and a value of 1 indicates that the medication was prescribed, while 0 indicates its absence. We will simplify the notation by omitting the superscript $j$ whenever there is no ambiguity.
@@ 4.3. Fine-tune and Inference
where $\hat{\mathbf{o}}\in[0,1]^{|\mathcal{M}|}$ is the probability of medications being included in the prediction set. $W_{3}\in\mathbb{R}^{|\mathcal{M}|\times dim}$ and $b_{3}\in\mathbb{R}^{|\mathcal{M}|}$ are learnable parameters.
@@
The balance between accuracy and safety in medication recommendation is achieved by combining the losses using a weighted sum:
@@ 5.1. Experimental Setup
Evaluation Protocol. We evaluate the overall performance of all methods using widely accepted metrics (Shang et al., 2019b; Yang et al., 2021b; Wu et al., 2022; Sun et al., 2022; Yang et al., 2023), including Jaccard coefficient (Jaccard), Precision-Recall Area Under Curve (PRAUC), F1-score (F1), DDI rate (DDI), and the average number of recommended medicines (#MED), providing a comprehensive assessment of overall recommendation performance. It is worth noting that higher values of Jaccard, PRAUC, and F1 indicate improved accuracy, while a lower DDI value suggests enhanced medication safety. Furthermore, we believe that a successful medication recommender should resemble the behavior of doctors by recommending a similar number of medications, as reflected in the metric #MED. The computation of the above metrics are provided in Appendix B.2.
To examine the fairness dimension, we conduct a group-wise comparison by categorizing all patients into five distinct groups, denoted as $\{\mathcal{G}_{1},\cdots,\mathcal{G}_{5}\}$ , according to the occurrence frequency of the least common disease codes they exhibit. $\mathcal{G}_{1}$ represents patients with the most commonly observed clinical codes, while $\mathcal{G}_{5}$ comprises patients with the rarest codes. We calculate the average Jaccard coefficients for each of these five groups to assess the accuracy of recommendations for patients with different code occurrences. Additionally, we compute the standard deviation (denoted as $\sigma$) of these five Jaccard coefficients to evaluate whether the models perform similarly across various groups, indicating fairness in the recommendations. We further provide a case study in Appendix C.1 to present specific examples and scenarios of patients with rare clinical codes.
Compared Methods. We compare RAREMed against (1) Instance-based methods: LR, LEAP (Zhang et al., 2017), and (2) Longitudinal-based methods: RETAIN (Choi et al., 2016), G-Bert (Shang et al., 2019a), GAMENet (Shang et al., 2019b), SafeDrug (Yang et al., 2021b), COGNet (Wu et al., 2022), MICRON (Yang et al., 2021a) and MoleRec (Yang et al., 2023). Details of these models are provided in Appendix B.1. In addition to the medication recommendation methods mentioned above, we also implement a baseline approach specifically designed for fair recommendation in group-wise comparison:
  • Rebalancing: Pre-processing the data to generate fair training data is an effective approach to obtain fair models (Kamiran and Calders, 2009; Ekstrand et al., 2018). We employ a resampling technique to address the issue of data imbalance. Specifically, inspired by the Inverse Propensity Score (IPS-based) approach (Schnabel et al., 2016; Yang et al., 2018), we calculate the IPS score for each disease code. Subsequently, we assign each patient a weighted score based on the largest IPS score associated with their disease codes (corresponding to the rarest disease). In each training epoch, we perform random sampling with replacement from the training data, ensuring an equal number of patients. The probability of selecting a particular patient is directly proportional to their assigned patient weight. This approach guarantees that patients with rarer clinical codes have a higher likelihood of being included in the training data, thus effectively mitigating the adverse impact of data imbalance on the fairness of the recommendation process. We conduct experiment for this approach on our RAREMed without pre-training.
Implementation Details. The hyper-parameters of RAREMed and all baseline models are determined based on their performance on the validation set or as specified in their respective papers. Our RAREMed is pre-trained on the training set in a sequential manner, first on the SMP task and then on the SP task, for 30 epochs respectively. The transformer encoder in our model consists of 3 layers with 4 attention heads, and the embedding dimension is set to 512. The weights of the loss function, denoted as $\alpha$ and $\beta$, are set to 0.03 and 0.7, respectively. The parameters are trained using the AdamW optimizer (Loshchilov and Hutter, 2017) with a learning rate of 1e-5 and weight decay of 0.1.
5.2. Overall Performance Comparison (RQ1)
  • Baseline Comparison. LR and RETAIN, which were not originally designed for medication recommendation tasks, exhibit poor performance. LR overlooks the interaction among clinical codes, while RETAIN fails to adequately capture the current patient condition. Similarly, LEAP and G-Bert, which do not consider the procedure sequence, also underperform in this domain. Furthermore, COGNet and MICRON utilize historical information in a specific manner, but their performance significantly declines in the single-visit setting. By integrating external knowledge such as EHR graph and drug molecule structures, GAMENet, SafeDrug and MoleRec achieve better performance. Interestingly, these three methods perform even worse in multi-visit setting compared to single-visit setting with MIMIC-III, highlighting the complexity of leveraging historical information. Further discussion on this topic can be found in Appendix C.2.
5.3. Group-wise Performance Comparison (RQ2)
  • Baseline Methods Yield Unfair Recommendations. Conventional medication recommendation methods exhibit a notable decrease in accuracy as the frequency of disease codes decreases, as illustrated in Figure 3(a). This leads to a higher standard deviation of Jaccard across different groups, as depicted in Figure 3(b). Notably, G-Bert, which utilizes pre-training techniques, struggles to achieve satisfactory accuracy for patients with rare diseases, potentially due to its disregard for procedure codes and subsequent failure to capture a comprehensive representation of patients’ clinical information during pre-training. This underscores the superiority of our model structure and pre-training task design. Additionally, Rebalancing, tailored to address fairness concerns, also underperforms compared to RAREMed in terms of both fairness and overall performance.
  • Pre-training Demonstrates Effectiveness on Biased Datasets. The results presented in Figure 3(c) and (d) highlight the overall positive impact of our pre-training strategy on both overall performance and group-wise accuracy, with few exceptions. Specifically, on MIMIC-III, where the disease distribution is less skewed, our pre-training strategy consistently enhances performance across various groups, as evidenced by the uniform improvement ratio observed after 30 pre-training epochs. In contrast, on MIMIC-IV, characterized by a more biased disease distribution and presenting greater challenges666 This is supported by the extensive distribution of long-tail disease codes in MIMIC-IV, as illustrated in Figure 1, and the greater variance of Jaccard scores across groups, as presented in Table 4, in comparison to the findings in Figure 3(b). , RAREMed notably improves the performance of the rarest group, effectively mitigating unfairness. The substantial improvement in overall accuracy on such a more biased dataset is also noteworthy.
Model Overall Performance Group-wise Performance
$w/o$ P 0.4561 0.6939 0.6092 0.0509 12.94 0.4719 0.4666 0.4534 0.4475 0.4410 0.01157
$w/o$ U 0.4589 0.6965 0.6122 0.0545 13.13 0.4773 0.4707 0.4558 0.4485 0.4423 0.01320
$w/o$ s&r 0.4528 0.6892 0.6062 0.0554 13.15 0.4695 0.4646 0.4508 0.4451 0.4342 0.01286
5.4. Ablation Study (RQ3)
  • $w/o$ P”, which removes two pre-training tasks, resulting in initializing parameters of encoder and embedding layers randomly.
  • $w/o$ U”, which disables the unified encoder and learns two encoders for diseases and procedures seperately and then concatenates the output. This setup retains two pre-training tasks.
  • $w/o$ s&r”, it disables segment and relevance embedding layers.
  • $\mathbf{w/o}$ P”. As expected, the lower performance of the ”$w/o$ P” variant confirms the significance of our pre-training tasks. The improvement of RAREMed can be attributed to the enhanced encoder obtained through the pre-training step, which enables RAREMed to comprehend intricate disease/procedure associations and generate more comprehensive patient representations.
5.5. Hyperparameter Studies (RQ4)
Loss Weight. Figure 4(a) demonstrates that the value of $\alpha$ significantly affects the size of the recommended medicine set. Additionally, we observed that the recommendation accuracy initially increases and then declines as the number of recommended medications increases. This finding suggests that, in the field of medication recommendation, it is important to recommend a set of medications rather than the typical approach of generating a ranked list, as the size of the recommended set is also a crucial factor to consider.
Number of Pre-training Epochs. Figure 4(c) illustrates the impact of the number of pre-training epochs on the pre-training tasks and the medication recommendation task. It is evident that as the number of pre-training epochs increases, the model’s performance in pre-training tasks improves. This observation indicates that the model acquires knowledge about clinical code associations and gains comprehensive patient representations during pre-training, which are crucial for medication recommendation, particularly for patients with rare diseases. However, the improvement in the downstream task starts to diminish after approximately 30 epochs. This phenomenon may be attributed to overfitting in the pre-training task and the inherent differences between the pre-training and downstream tasks. These findings highlight the significance of striking an optimal balance between the performance on pre-training tasks and the downstream medication recommendation task.
Figure 4. Hyperparameter effects on model performance.
6. Conclusion
In this paper, we propose RAREMed, a novel medication recommendation model designed to addresses the fairness issue in medication recommendation systems. By focusing on enhancing accuracy for patients with rare diseases, RAREMed utilizes a unified input sequence approach to capture complex relationships among disease and procedure codes. The model also leverages self-supervised pre-training tasks to learn specialized medication needs and interrelations. Experimental results on real-world datasets demonstrate the effectiveness of RAREMed in providing accurate drug recommendations for both rare and common disease patients, thereby mitigating unfairness.
Acknowledgements.
We thank the anonymous reviewers for taking time to read and make valuable comments on this paper.
Appendix A Profile of Rare Disease Patients
In this section, we conduct empirical analysis on real-world EHR dataset MIMIC-IV (Johnson et al., 2023) to explore the profile of patients with rare diseases. Specifically, we categorize patients into 13 groups based on the popularity of their rarest disease codes, using uniform segmentation across popularity intervals. We then calculate various metrics such as average disease count, procedure count, medication count, and medication popularity for each group. From the results presented in Figure 5, we make two important observations:
Observation 1: The data presented in Figure 5(a) and 5(b) clearly demonstrates a negative correlation between disease count and procedure count with patient disease popularity. This suggests that as the rarity of a patient’s disease increases, their clinical condition becomes more complex. Consequently, patients with rare diseases exhibit more intricate combinations of clinical conditions.
Observation 2: Figure 5(c) and 5(d) demonstrate that patients with rarer diseases tend to have a higher number of prescribed medications and a lower popularity of these medications. This indicates that rare diseases patients require more specifically tailored medication options, resulting in more complex therapeutic regimens.
Appendix B Additional Experiment Details
B.1. Baseline Details
@@
We compare RAREMed against the following baseline algorithms:

### v2

@@ Abstract.
Medication recommendation systems have gained significant attention in healthcare as a means of providing tailored and effective drug combinations based on patients’ clinical information. However, existing approaches often suffer from fairness issues, as recommendations tend to be more accurate for patients with common diseases compared to those with rare conditions. In this paper, we propose a novel model called Robust and Accurate REcommendations for Medication (RAREMed), which leverages the pretrain-finetune learning paradigm to enhance accuracy for rare diseases. RAREMed employs a transformer encoder with a unified input sequence approach to capture complex relationships among disease and procedure codes. Additionally, it introduces two self-supervised pre-training tasks, namely Sequence Matching Prediction (SMP) and Self Reconstruction (SR), to learn specialized medication needs and interrelations among clinical codes. Experimental results on two real-world datasets demonstrate that RAREMed provides accurate drug sets for both rare and common disease patients, thereby mitigating unfairness in medication recommendation systems. The implementation is available via https://github.com/zzhUSTC2016/RAREMed
Medication recommendation; Electronic health record; Fairness; Rare disease
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USA.booktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAisbn: 979-8-4007-0431-4/24/07doi: 10.1145/3626772.3657785ccs: Information systems Recommender systemsccs: Applied computing Health informatics
@@ 1. Introduction
Current approaches to medication recommendation primarily focus on maximizing overall accuracy by incorporating patients’ current medical conditions (Zhang et al., 2017; Wang et al., 2019) or clinical information from previous visits (Shang et al., 2019a, b; Yang et al., 2021b; Wu et al., 2022; Yang et al., 2021a; Sun et al., 2022). While significant progress has been made in comprehensive utilization these data, existing approaches encounter fairness issues in terms of recommendations across patient groups. This stems from the highly skewed distribution of clinical codes in Electronic Health Records (EHR) (Kim et al., 2019) — a digital compilation of patients’ medical conditions, including diseases, procedures, medications and other clinical data during patient visits. As illustrated in Figure 1(a), a small subset of codes exhibit high prevalence, while the long tail comprises low-occurrence diseases in two real-world EHR datasets. Consequently, patients resided in the most common group receive significantly more accurate recommendations compared to those in the rarest group, as depicted in Figure 1(b). This disparity in accuracy across patient groups raises concerns regarding equity and fairness, ultimately undermining the reliability and utility of medication recommendation systems.
@@
  • Tailored medication needs [^1] We will further discuss the profile of rare disease patients in Section 4.: The atypical pathophysiology of rare diseases often necessitates more precisely targeted pharmacological interventions. On the other hand, patients with rare diseases tend to have more complicated therapeutic regimens, warrant tailored medication recommendations.
@@ 2. Related Works
Medication Recommendation. Methods for medication recommendation can be broadly categorized into two groups: longitudinal and instance-based approaches. Longitudinal methods make use of patients’ longitudinal medical history, such as the approach taken by Choi et al. (Choi et al., 2016), who employ a two-level temporal attention mechanism. Similarly, Shang et al. (Shang et al., 2019a) pre-train their model on single-visit data [^3] Single-visit patients have no historical visits, i.e., have only one visit recorded in EHR. and fine-tune it using multi-visit data [^4] Multi-visit patients have at least two visits in EHR.. Additionally, Shang et al. (Shang et al., 2019b) incorporate a graph-augmented memory module and DDI graph to reduce adverse drug-drug interactions, while Yang et al. (Yang et al., 2021b) consider drug molecule structures for medication security. Wu et al. (2022) introduce a copy-or-predict mechanism, and Yang et al. (Yang et al., 2021a) propose a residual-based recurrent network. Sun et al. (Sun et al., 2022) present a causal model to address recommendation bias, and Yang et al. (Yang et al., 2023) leverage a molecular substructure-aware attentive method. More recently, Bhoi et al. (Bhoi et al., 2023) propose a fine-grained medication recommendation approach. However, many of these methods heavily rely on historical records, making them less suitable for single-visit patients (Choi et al., 2016; Shang et al., 2019a; Sun et al., 2022; Bhoi et al., 2023), or exhibit reduced accuracy in such cases (Shang et al., 2019b; Yang et al., 2021b; Wu et al., 2022; Yang et al., 2021a, 2023). Instance-based methods focus on the current medical condition of patients. For instance, Zhang et al. (Zhang et al., 2017) utilize an RNN model based on disease codes. Wang et al. (Wang et al., 2019) employ Deep Q Learning to capture correlations and adverse interactions between medicines. However, these models may not adequately capture patient-specific information, potentially leading to reduced accuracy. Despite the initial success of these methods, a common issue among them is fairness (Chen et al., 2023; Wang et al., 2023). Patients with common diseases tend to receive more accurate medication predictions compared to those with rare diseases. This discrepancy undermines the overall performance and impairs the reliability and usefulness of medication recommendation systems.
@@ 3. Problem Formulation
The record of patient $j$, denoted as $\mathcal{V}^{(j)}=\{\mathbf{d}^{(j)},\mathbf{p}^{(j)},\mathbf{m}^{(j)}\}$, comprises three components. Firstly, $\mathbf{d}^{(j)}=[d_{1},d_{2},\cdots,d_{x}]\in\mathcal{D}$ represents the sequence of diseases experienced by the patient, ordered by their priority or significance [^5] The sequence numbers of priority are provided in the EHR datasets (Johnson et al., 2016, 2023), which are labeled by qualified physicians., reflecting their relative importance in the patient’s healthcare journey. Similarly, $\mathbf{p}^{(j)}=[p_{1},p_{2},\cdots,p_{y}]\in\mathcal{P}$ captures the sequence of procedures, also ordered by priority. Here, $x$ and $y$ denote the lengths of the disease and procedure sequences, respectively. Lastly, $\mathbf{m}^{(j)}\in\{0,1\}^{|\mathcal{M}|}$ is a multi-hot vector indicating the medications prescribed to the patient. Each element in $\mathbf{m}^{(j)}$ corresponds to a specific medication in $\mathcal{M}$, and a value of 1 indicates that the medication was prescribed, while 0 indicates its absence. We will simplify the notation by omitting the superscript $j$ whenever there is no ambiguity.
@@
4. Profile of Rare Disease Patients
In this section, we conduct empirical analysis on real-world EHR dataset MIMIC-IV (Johnson et al., 2023) to explore the profile of patients with rare diseases. Specifically, we categorize patients into 13 groups based on the popularity of their rarest disease codes, using uniform segmentation across popularity intervals. We then calculate various metrics for each group. From the results presented in Figure 2, we make two important observations:
Observation 1: The data presented in Figure 2(a) and 2(b) clearly demonstrates a negative correlation between disease count and procedure count with patient disease popularity. This suggests that as the rarity of a patient’s disease increases, their clinical condition becomes more complex. Consequently, patients with rare diseases exhibit more intricate combinations of clinical conditions.
Observation 2: Figure 2(c) and 2(d) demonstrate that patients with rarer diseases tend to have a higher number of prescribed medications and a lower popularity of these medications. This indicates that rare diseases patients require more specifically tailored medication options, resulting in more complex therapeutic regimens.
@@ 5.3. Fine-tune and Inference
where $\hat{\mathbf{o}}\in[0,1]^{|\mathcal{M}|}$ is the probability of medications being recommended. $W_{3}\in\mathbb{R}^{|\mathcal{M}|\times dim}$ and $b_{3}\in\mathbb{R}^{|\mathcal{M}|}$ are learnable parameters.
@@
The balance between accuracy and safety is achieved by combining the losses using a weighted sum:
@@ 6.1. Experimental Setup
Evaluation Protocol. We evaluate the overall performance of all methods using widely accepted metrics (Shang et al., 2019b; Yang et al., 2021b; Wu et al., 2022; Sun et al., 2022; Yang et al., 2023), including Jaccard coefficient (Jaccard), Precision-Recall Area Under Curve (PRAUC), F1-score (F1), DDI rate (DDI), and the average number of recommended medicines (#MED), providing a comprehensive assessment of overall recommendation performance. It is worth noting that higher values of Jaccard, PRAUC, and F1 indicate improved accuracy, while a lower DDI value suggests enhanced medication safety. Furthermore, we believe that a successful medication recommender should resemble the behavior of doctors by recommending a similar number of medications, as reflected in the metric #MED.
Details of metrics are listed below. Building upon the formulation detailed in Section 3, $\mathcal{M}$ represents the set of all medications in the dataset, while $\mathbf{m},\hat{\mathbf{m}}\in\{0,1\}^{|\mathcal{M}|}$ signify the prescribed and predicted medications to a patient, respectively.
$\text{F}_{1}=\frac{2\text{R}\times\text{P}}{\text{R}+\text{P}},$
where the recall and precision are formulated as
where $\mathbf{A}$ is the DDI graph defined in Section 3.
To examine the fairness dimension, we conduct a group-wise comparison by categorizing all patients into five groups, denoted as $\{\mathcal{G}_{1},\cdots,\mathcal{G}_{5}\}$, according to the occurrence frequency of the least common disease codes they exhibit. $\mathcal{G}_{1}$ represents patients with the most commonly observed clinical codes, while $\mathcal{G}_{5}$ comprises patients with the rarest codes. We calculate the average Jaccard coefficients for each group to assess the accuracy of recommendations for patients with different code occurrences. Additionally, we compute the standard deviation (denoted as $\sigma$) of these five Jaccard coefficients to evaluate whether the models perform similarly across various groups, indicating fairness in the recommendations.
@@
Compared Methods. We compare RAREMed against the following baseline algorithms: